# Radial Oscillations of Viscous Neutron Stars: Zero Diffusion Case

### v1

@@ IV.1 Zero energy diffusion limit
Physical solutions for the Lagrangian displacement must be regular throughout the domain, which imposes the boundary condition $\xi(r=0)=0$ and a Robin-type boundary condition (i.e., $a\xi^{\prime}(R)+b\xi(R)=0$ for some $a,b\in\mathbb{C}$) at the stellar surface ($r=R$), the precise form of which depends on the choice of transport coefficients. Since the master equation for $\xi$ in the frequency domain is linear and homogeneous, it admits a rescaling freedom: if $\xi$ is a solution, then so is $c\xi$, for any $c\in\mathbb{C}$. Therefore, the boundary conditions at $r=0$ and $r=R$ cannot be simultaneously satisfied for arbitrary values of $\omega$, but only for a (possibly infinite) discrete set of eigenfrequencies, which we denote by $\omega_{(n)}$. Here, the index $n\in\mathbb{N}$ labels the eigenmodes and corresponds to the number of nodes (zeros) in the radial profile of the eigenfunctions.
@@ VI Conclusions
For $n\geq 1$ overtones, the effects of bulk and shear viscosity are of comparable magnitude. However, shear viscosity is shown to leave the fundamental ($n=0$) modes largely unchanged, whereas bulk viscosity strongly affects them (cf. Fig. 2) – a behavior consistent with the symmetry of the problem. The influence of bulk viscosity on the $n=0$ modes is also shown to be highly sensitive to relativistic corrections. This is illustrated for Navier-Stokes theory in Fig. 3(b) and for Israel-Stewart theory in Figs. 4(b) and 7, the latter employing a more realistic EOS. In Israel-Stewart theory, we further show that, depending on the strength of bulk-viscous effects, stability against gravitational collapse may be determined by the $n=0$ nonhydrodynamic mode rather than the hydrodynamic one. The strikingly different impacts of shear and bulk viscosity can also be visualized in the time-domain simulation shown in Fig. 5.
@@ Appendix B Longitudinal perturbations around a homogeneous fluid in a flat spacetime
The discussion in this section summarizes what is presented in Ref. [46, cap. 2]. We consider a fluid initially in global equilibrium, such that its state is specified by $\{\mathrm{n}_{(0)},\mathrm{e}_{(0)},\mathrm{p}_{(0)}(\mathrm{n}_{(0)},\mathrm{e}_{(0)}),u^{\mu}_{(0)}\}$. Then, it is linearly perturbed to a non-equilibrium state specified by $(\mathrm{n}_{(0)}+\bar{\delta}\mathrm{n},\mathrm{e}_{(0)}+\bar{\delta}\mathrm{e},\mathrm{p}_{(0)}(\mathrm{n}_{(0)},\mathrm{e}_{(0)})+\bar{\delta}\mathrm{p}_{(0)}+\bar{\Pi},u^{\mu}_{(0)}+\bar{\delta}u^{\mu},\bar{\pi}^{\alpha\beta})$. Then, from the linearized local conservation laws (7) in Fourier space, in the limit where particle density and energy diffusion perturbations vanish, we have
@@
For Navier-Stokes in Fourier space we have
$\bar{\Pi}=-\zeta i\kappa_{\mu}\bar{\delta}u^{\mu}$, $\bar{\pi}^{\alpha\beta}=-2i\eta\Delta^{\alpha\beta\mu\nu}_{(0)}\kappa_{\mu}\bar{\delta}u_{\nu}$, where $\Delta^{\alpha\beta\mu\nu}_{(0)}$is the doubly-symmetric, traceless projector that is fully orthogonal with respect to the background four-velocity. Hence, Eqs. (74) can be expressed as
$\left(\begin{array}[]{cc}-i\Omega&i\kappa\\ i\kappa c_{s}^{2}&-i\Omega-2\kappa^{2}/k_{\rm visc}\end{array}\right)\left(\begin{array}[]{c}\bar{\delta}\mathrm{e}/(\mathrm{e}+\mathrm{p})\\ \bar{\delta}u_{\parallel}\end{array}\right)=\left(\begin{array}[]{c}0\\ 0\end{array}\right)$ (76)
@@
$\Omega^{2}-2ic_{s}\frac{\kappa^{2}}{k_{\rm visc}}\Omega-\kappa^{2}c_{s}^{2}=0$ (77)
which in the local rest frame of the background fluid can be solved for the frequency and we obtain the dispersion relation (48).
On the other hand, for Israel Stewart we have, in Fourier space $(-i\Omega+1)\bar{\Pi}=-\zeta\kappa_{\mu}\bar{\delta}u^{\mu}$, for the bulk relaxation equation and $(-i\Omega+1)\bar{\pi}^{\alpha\beta}=-2i\eta\Delta^{\alpha\beta\mu\nu}_{(0)}\kappa_{\mu}\bar{\delta}u_{\nu}$ for the shear relaxation equation. For systems, with only bulk dissipative perturbation, Eqs. (74) can be expressed as
$\left(\begin{array}[]{ccc}-i\Omega&i\kappa&0\\ i\kappa c_{s}^{2}&-i\Omega&i\kappa\\ 0&2i\kappa/k_{\rm visc}|_{\eta=0}&(-i\Omega\tau_{\Pi}+1)\end{array}\right)\left(\begin{array}[]{c}\bar{\delta}\mathrm{e}/(\mathrm{e}+\mathrm{p})\\ \bar{\delta}u_{\parallel}\\ \bar{\Pi}/(\mathrm{e}+\mathrm{p})\end{array}\right)$ (78)
@@
$(\Omega^{2}-c_{s}^{2}\kappa^{2})(1-i\Omega\tau_{\pi})+2i\Omega c_{s}\kappa\left(\frac{\kappa}{k_{\textrm{visc}}|_{\zeta=0}}\right)=0.$ (79)
which in the local rest frame of the background fluid reduces to Eq. (51). Alternatively, for systems with only shear dissipative perturbations, Eqs. (74) can be expressed as
$\left(\begin{array}[]{ccc}-i\Omega&i\kappa&0\\ i\kappa c_{s}^{2}&-i\Omega&i\kappa\\ 0&2i\kappa/k_{\rm visc}|_{\zeta=0}&(-i\Omega\tau_{\pi}+1)\end{array}\right)\left(\begin{array}[]{c}\bar{\delta}\mathrm{e}/(\mathrm{e}+\mathrm{p})\\ \bar{\delta}u_{\parallel}\\ \bar{\pi}_{\parallel}/(\mathrm{e}+\mathrm{p})\end{array}\right)$ (80)
@@
which is just Eq. (78), with $\tau_{\Pi}\mapsto\tau_{\pi}$, $k_{\rm visc}|_{\eta=0}\mapsto k_{\rm visc}|_{\zeta=0}$ and $\bar{\Pi}\mapsto\bar{\pi}_{\parallel}$. Thus, imposing the vanishing of the above-defined matrix determinant amounts to Eq. (79) with $\tau_{\Pi}\mapsto\tau_{\pi}$, $k_{\rm visc}|_{\eta=0}\mapsto k_{\rm visc}|_{\zeta=0}$,
$(\Omega^{2}-c_{s}^{2}\kappa^{2})(1-i\Omega\tau_{\Pi})+2i\Omega c_{s}\kappa\left(\frac{\kappa}{k_{\textrm{visc}}|_{\eta=0}}\right)=0.$ (81)

### v2

@@ IV.1 Zero energy diffusion limit
Physical solutions for the Lagrangian displacement must be regular throughout the domain, which imposes the boundary condition $\xi(r=0)=0$ and a Robin-type boundary condition (i.e., $a\xi^{\prime}(R)+b\xi(R)=0$ for some $a,b\in\mathbb{C}$) at the stellar surface ($r=R$), the precise form of which depends on the choice of transport coefficients. Since in the frequency domain bulk- and shear-stress perturbations are algebraically related to the Lagrangian displacement and its derivative, $\Pi(R)$ and $\tilde{\pi}(R)$ are completely determined by $\xi(R)$ and $\xi^{\prime}(R)$; no additional boundary conditions for the viscous variables are needed at the stellar surface.
Since the master equation for $\xi$ in the frequency domain is linear and homogeneous, it admits a rescaling freedom: if $\xi$ is a solution, then so is $c\xi$, for any $c\in\mathbb{C}$. Therefore, the boundary conditions at $r=0$ and $r=R$ cannot be simultaneously satisfied for arbitrary values of $\omega$, but only for a (possibly infinite) discrete set of eigenfrequencies, which we denote by $\omega_{(n)}$. Here, the index $n\in\mathbb{N}$ labels the eigenmodes and corresponds to the number of nodes (zeros) in the radial profile of the eigenfunctions.
@@ VI Conclusions
For $n\geq 1$ overtones, the effects of bulk and shear viscosity are of comparable magnitude. However, shear viscosity is shown to leave the fundamental ($n=0$) modes largely unchanged, whereas bulk viscosity strongly affects them (cf. Fig. 2) – a behavior consistent with the symmetry of the problem. The influence of bulk viscosity on the $n=0$ modes is also shown to be highly sensitive to relativistic corrections. This is illustrated for Navier-Stokes theory in Fig. 3(b) and for Israel-Stewart theory in Figs. 4(b) and 7, the latter employing a more realistic EOS. In Israel-Stewart theory, we further show that, depending on the strength of bulk-viscous effects, stability against gravitational collapse may be determined by the $n=0$ nonhydrodynamic mode rather than the hydrodynamic one. The strikingly different impacts of shear and bulk viscosity can also be visualized in the time-domain simulation shown in Fig. 5. As far as unstable radial modes are concerned, we find that viscosity acts to soften the instability (i.e., it increases the growth timescale of the perturbations); however, it is not able to suppress it altogether, consistent with the conclusion from the energy-based stability analysis in Ref. [37].
@@ Appendix B Longitudinal perturbations around a homogeneous fluid in a flat spacetime
The discussion in this section summarizes what is presented in Ref. [46, cap. 2]. We consider a fluid initially in global equilibrium, such that its state is specified by $\{\mathrm{n}_{(0)},\mathrm{e}_{(0)},\mathrm{p}_{(0)}(\mathrm{n}_{(0)},\mathrm{e}_{(0)}),u^{\mu}_{(0)}\}$. Then, it is linearly perturbed to a non-equilibrium state specified by $(\mathrm{n}_{(0)}+\bar{\delta}\mathrm{n},\mathrm{e}_{(0)}+\bar{\delta}\mathrm{e},\mathrm{p}_{(0)}(\mathrm{n}_{(0)},\mathrm{e}_{(0)})+\bar{\delta}\mathrm{p}_{(0)}+\bar{\Pi},u^{\mu}_{(0)}+\bar{\delta}u^{\mu},\bar{\pi}^{\alpha\beta})$. Then, from the linearized local conservation laws (7) in Fourier space, in a regime where energy diffusion perturbations vanish, we have
@@
For Navier-Stokes in Fourier space we have $\bar{\Pi}=-\zeta i\kappa_{\mu}\bar{\delta}u^{\mu}$, $\bar{\pi}^{\alpha\beta}=-2i\eta\Delta^{\alpha\beta\mu\nu}_{(0)}\kappa_{\mu}\bar{\delta}u_{\nu}$, where $\Delta^{\alpha\beta\mu\nu}_{(0)}$ is the doubly-symmetric, traceless projector that is fully orthogonal with respect to the background four-velocity. Hence, Eqs. (74) can be expressed as
$\left(\begin{array}[]{cc}-i\Omega&i\kappa\\ i\kappa c_{s}^{2}&-i\Omega+2c_{s}\kappa^{2}/k_{\rm visc}\end{array}\right)\left(\begin{array}[]{c}\bar{\delta}\mathrm{e}/(\mathrm{e}+\mathrm{p})\\ \bar{\delta}u_{\parallel}\end{array}\right)=\left(\begin{array}[]{c}0\\ 0\end{array}\right).$ (76)
@@
$\Omega^{2}+2ic_{s}\frac{\kappa^{2}}{k_{\rm visc}}\Omega-\kappa^{2}c_{s}^{2}=0,$ (77)
which in the local rest frame of the background fluid can be solved for the frequency and yields the dispersion relation (48).
On the other hand, for Israel-Stewart we have, in Fourier space $(1-i\Omega\tau_{\Pi})\bar{\Pi}=-\zeta i\kappa_{\mu}\bar{\delta}u^{\mu}$, for the bulk relaxation equation and $(1-i\Omega\tau_{\pi})\bar{\pi}^{\alpha\beta}=-2i\eta\Delta^{\alpha\beta\mu\nu}_{(0)}\kappa_{\mu}\bar{\delta}u_{\nu}$ for the shear relaxation equation. For systems with only bulk dissipative perturbations, Eqs. (74) can be expressed as
$\left(\begin{array}[]{ccc}-i\Omega&i\kappa&0\\ i\kappa c_{s}^{2}&-i\Omega&i\kappa\\ 0&2ic_{s}\kappa/k_{\rm visc}|_{\eta=0}&(-i\Omega\tau_{\Pi}+1)\end{array}\right)\left(\begin{array}[]{c}\bar{\delta}\mathrm{e}/(\mathrm{e}+\mathrm{p})\\ \bar{\delta}u_{\parallel}\\ \bar{\Pi}/(\mathrm{e}+\mathrm{p})\end{array}\right)$ (78)
@@
$(\Omega^{2}-c_{s}^{2}\kappa^{2})(1-i\Omega\tau_{\Pi})+2i\Omega c_{s}\kappa\left(\frac{\kappa}{k_{\textrm{visc}}|_{\eta=0}}\right)=0,$ (79)
which in the local rest frame of the background fluid reduces to Eq. (51). Conversely, for systems with only shear dissipative perturbations, Eqs. (74) can be expressed as
$\left(\begin{array}[]{ccc}-i\Omega&i\kappa&0\\ i\kappa c_{s}^{2}&-i\Omega&i\kappa\\ 0&2ic_{s}\kappa/k_{\rm visc}|_{\zeta=0}&(-i\Omega\tau_{\pi}+1)\end{array}\right)\left(\begin{array}[]{c}\bar{\delta}\mathrm{e}/(\mathrm{e}+\mathrm{p})\\ \bar{\delta}u_{\parallel}\\ \bar{\pi}_{\parallel}/(\mathrm{e}+\mathrm{p})\end{array}\right)$ (80)
@@
which is just Eq. (78), with $\tau_{\Pi}\mapsto\tau_{\pi}$, $k_{\rm visc}|_{\eta=0}\mapsto k_{\rm visc}|_{\zeta=0}$ and $\bar{\Pi}\mapsto\bar{\pi}_{\parallel}$. Thus, imposing that the determinant of the above-defined matrix vanishes amounts to Eq. (79) with $\tau_{\Pi}\mapsto\tau_{\pi}$, $k_{\rm visc}|_{\eta=0}\mapsto k_{\rm visc}|_{\zeta=0}$,
$(\Omega^{2}-c_{s}^{2}\kappa^{2})(1-i\Omega\tau_{\pi})+2i\Omega c_{s}\kappa\left(\frac{\kappa}{k_{\textrm{visc}}|_{\zeta=0}}\right)=0,$ (81)